OpenStudy (anonymous):

Vertex form to standard f(x) = (x - 4)^2 -4

OpenStudy (xapproachesinfinity):

Just expand (x-4)^2 do you know how to do this?

OpenStudy (anonymous):

it becomes (x-2)(x-2)

OpenStudy (xapproachesinfinity):

no! it is (x-4)(x-4) then FOIL

OpenStudy (anonymous):

thats where I mess up.

OpenStudy (xapproachesinfinity):

\(\Large \tt\color{blueviolet}{(x-4)(x-4)=x^2-4x-4x+16=x^2-8x+16}\)

OpenStudy (anonymous):

thank you so muchh!!

OpenStudy (xapproachesinfinity):

yw! you are not finished yet you know

OpenStudy (nincompoop):

so what happened to the -4?

OpenStudy (anonymous):

u add +4??

OpenStudy (xapproachesinfinity):

no you subtract 4 from what we got above

OpenStudy (xapproachesinfinity):

\(\Large \tt\color{blueviolet}{f(x)=x^2-8x+16-4}\)

OpenStudy (xapproachesinfinity):

now complete the it to the final answer

OpenStudy (anonymous):

f(x)=x^2-8x+12

OpenStudy (xapproachesinfinity):

Correct!

OpenStudy (anonymous):

thank youu

OpenStudy (xapproachesinfinity):

Anytime!

OpenStudy (nincompoop):

hmm

OpenStudy (xapproachesinfinity):

@nincompoop hehe

OpenStudy (nincompoop):

f(x) = a(x-h)^2 + k try to convert f(x) = 2(x+2)^2 - 1 into standard-form just to make sure you learned what you were taught

OpenStudy (xapproachesinfinity):

\(\Large \tt\color{blueviolet}{He~is~gone}\) not sure he got this lol